help me with the problem below

my percent body fat is 95%
I weight 160 pounds
my height is 5"2



There are several different ways to compute an ideal body weight. Method A of this laboratory enables you to compute and record your ideal body weight using skinfold measurements. (In Chapter 6 we discussed body fat percentage estimated from skinfold measurements.) Method B enables you to calculate and record your ideal body weight using the body mass index (BMI) procedure (Chapter 6). Choose one of these techniques, and complete the appropriate section.


METHOD A: COMPUTING IDEAL BODY WEIGHT USING PERCENT BODY FAT
STEP 1: Calculate fat-free weight
100% - your percent body fat estimated from skinfold measurement = , fat-free weight
Therefore,


, fat-free weight expressed as a decimal * your body weight in pounds Calculate
= pounds of fat-free weight

STEP 2: Calculate optimal weight
Remember: Optimal body fat ranges are 8%–19% for men and 21%–32% for women. Optimal weight = fat-free weight , (1.00 -
optimal % fat), with optimal % fat expressed as a decimal. Therefore, the low and high optimal weight ranges for your gender are as follows:
For low ,fat : Optimal weight pounds

For high ,fat : Optimal weight pounds


METHOD B: COMPUTING IDEAL BODY WEIGHT USING BODY MASS INDEX (BMI)
The BMI uses the metric system. Therefore, you must express your weight in kilograms (1 kilogram = 2.2 pounds) and your height in meters (1 inch = 0.0254 meter).

STEP 1: Compute your BMI
BMI = body weight 1 kg 2 , 1 height in meters 2 2

Your BMI =

STEP 2: Calculate your ideal body weight based on BMI
The ideal BMI is 21.9 to 22.4 for men and 21.3 to 22.1 for women. The formula for computing ideal body weight using BMI is
Ideal body weight 1 kilograms 2 = Desired BMI * 1 height in meters 2 2
Consider the following example as an illustration of the computation of ideal body weight. A man who weighs 60 kilograms and is 1.5 meters tall computes his BMI to be 26.7. His ideal BMI is between 21.9 and 22.4; therefore, his ideal body weight range is as follows:
Low end range : 21.9 * 2.25 = 49.3 kilograms
High end of range: 22.4 * 2.25 = 50.4 kilograms
Now complete this calculation using your values for BMI.
My ideal body weight range using the BMI method is to kilograms.


Note: BMI may not be a good method to determine ideal body weight for a highly muscled individual.

what type of information you need?
Answer #1

BMI=WEIGHT IN KG/HEIGHT IN M2

BMI=72/(1.57)2=29 it is considered over weight.

Here not mention you are male ior female, Ideal body weight In male =21.9*2.46=54kg (low end range)

High range ,22.4*2.46=55kg